OpenStudy (anonymous):

evaluate following expression in simplest form lne^-1 eln^7 e^lnsqrt3 ln(1/e^5)

OpenStudy (anonymous):

note that \[e^x\] and \[ln(x)\] are inverse functions, so that \[\ln(e^x)=x,\text{ and } e^{\ln(x)}=x\]

OpenStudy (anonymous):

so you should just be able to write down the answers to the first three without thought

OpenStudy (anonymous):

got it